# Extended mass and spheroidal deformation effects on epicyclic frequencies and periapsis shift in quasi-circular orbits

### v2

@@ V.4 $n=3$
The upper panels of Fig. 7 show $\Delta\phi_{\mathrm{p}}$ and $f/(1+3\rho/\bar{\rho})$. Note, however, that marginally stable circular orbits do not appear, and quasi-circular orbits exist throughout the entire range $R\leq a$.
@@ V.5 $n=4$
The lower panels of Fig. 7 show $\Delta\phi_{\mathrm{p}}$ and $f/(1+3\rho/\bar{\rho})$. No marginally stable circular orbit appears, and quasi-circular orbits exist throughout the entire range $R\leq a$.
VI Summary and discussion
@@
We also analyzed the stability of quasi-circular orbits in this system by examining the epicyclic frequencies. In the uniform-density case, the quasi-circular orbits remain stable regardless of the guiding orbit radius or the deformation parameter. In contrast, when the density is inhomogeneous—especially under oblate deformation—the radial epicyclic frequency becomes zero, indicating that marginally stable circular orbits appear for $e_{0}\leq e<1$. This phenomenon seems to occur only for $n=1$ and $n=2$. Specifically, for $n=1$, the critical value is $e_{0}=0.929955685$, which agrees with the oblateness at which marginally stable circular orbits emerge in the external vacuum region [47]. This agreement occurs because, in both the interior region and the external vacuum region, the marginally stable circular orbit emerges at the spheroid’s surface when $e=e_{0}$. In contrast, for $n=2$, the critical value is $e_{0}=0.988937349$, and no such correspondence with the external vacuum region’s oblateness is observed. Notably, the marginally stable orbits—often misinterpreted as a phenomenon unique to general relativity—emerge under Newtonian conditions due to extended mass effects and density inhomogeneity. Meanwhile, under prolate deformation, marginally stable orbits are not formed.
@@
We will report the results of our post-Newtonian analysis in a separate paper, which examines the competition between the general-relativistic effects and the combined extended mass and deformation effects on the periapsis shift of quasi-circular orbits in a system featuring a black hole at the center of a spheroidal mass distribution. Furthermore, the extended mass and deformation effects on the periapsis shift of quasi-circular orbits in static, axisymmetric spacetimes warrant future work [51]. In addition, generalizing the general formula for the periapsis shift of quasi-circular orbits in static, spherically symmetric systems (2) to static, axisymmetric spacetimes is a promising topic for future research.
Acknowledgements.
@@
The author gratefully acknowledges the useful comments and discussions provided by Kota Ogasawara, Tomohiro Harada, Masashi Kimura, Keiju Murata, Takeshi Chiba, Masaya Amo, Akihito Katsumata, Takashi Mishima, Yohsuke Takamori, Takaaki Ishii, Hideki Ishihara, Ken-ichi Nakao, Yasufumit Kojima, Masaaki Takahashi, Rohta Takahashi, Shinji Koide, Hirotaka Yoshino, and So Aina. This work was supported by JSPS KAKENHI Grant Nos. JP22K03611, JP23KK0048, and JP24H00183.
@@ A.1 Oblate case ($a>b$)
$C_{5,0}$ $=\frac{63(1-e^{2})}{128e^{10}}\left(\frac{\arcsin e}{e\sqrt{1-e^{2}}}-1-\frac{2e^{2}}{3}-\frac{8e^{4}}{15}-\frac{16e^{6}}{35}-\frac{128e^{8}}{315}\right).$ (76)
@@
$C_{3,1}$ $=\frac{35(1-e^{2})}{8e^{8}}\left(1-\frac{e^{2}}{3}-\frac{2e^{4}}{15}-\frac{8e^{6}}{105}-\frac{\sqrt{1-e^{2}}}{e}\arcsin e\right),$ (80)
$C_{4,1}$ $=\frac{315(1-e^{2})}{64e^{10}}\left(1-\frac{e^{2}}{3}-\frac{2e^{4}}{15}-\frac{8e^{6}}{105}-\frac{16e^{8}}{315}-\frac{\sqrt{1-e^{2}}}{e}\arcsin e\right).$ (81)
@@ A.2 Prolate case ($a<b$)
$C_{4,0}$ $=\frac{35}{64e^{8}}\left(-1+\frac{11e^{2}}{3}-\frac{73e^{4}}{15}+\frac{93e^{6}}{35}+\frac{(1-e^{2})^{4}}{e}\operatorname{arctanh}e\right),$ (86)
$C_{5,0}$ $=\frac{63}{128e^{10}}\left(1-\frac{14e^{2}}{3}+\frac{128e^{4}}{15}-\frac{158e^{6}}{21}+\frac{193e^{8}}{63}-\frac{(1-e^{2})^{5}}{e}\operatorname{arctanh}e\right),$ (87)
@@
$C_{4,1}$ $=\frac{315}{64e^{10}}\left(-1+\frac{11e^{2}}{3}-\frac{73e^{4}}{15}+\frac{93e^{6}}{35}-\frac{138e^{8}}{315}+\frac{(1-e^{2})^{4}}{e}\operatorname{arctanh}e\right).$ (92)

### v3

@@ 5.4 $n=3$
The upper panels of Fig. 8 show $\Delta\phi_{\mathrm{p}}$ and $f/(1+3\rho/\bar{\rho})$. Note, however, that MSCOs do not appear, and quasi-circular orbits exist throughout the entire range $R\leq a$.
@@ 5.5 $n=4$
The lower panels of Fig. 8 show $\Delta\phi_{\mathrm{p}}$ and $f/(1+3\rho/\bar{\rho})$. No MSCO appears, and quasi-circular orbits exist throughout the entire range $R\leq a$.
6 Summary and discussion
@@
We also analyzed the stability of quasi-circular orbits in this system by examining the epicyclic frequencies. In the uniform-density case, the quasi-circular orbits remain stable regardless of the guiding orbit radius or the deformation parameter. In contrast, when the density is inhomogeneous—especially under oblate deformation—the radial epicyclic frequency becomes zero, indicating that MSCOs appear for $e_{0}\leq e<1$. This phenomenon seems to occur only for $n=1$ and $n=2$. Specifically, for $n=1$, the critical value is $e_{0}=0.929955685$, which agrees with the oblateness at which MSCOs emerge in the external vacuum region (Bollimpalli, 2022). This agreement occurs because, in both the interior region and the external vacuum region, the MSCO emerges at the spheroid’s surface when $e=e_{0}$. In contrast, for $n=2$, the critical value is $e_{0}=0.988937349$, and no such correspondence with the external vacuum region’s oblateness is observed. Notably, the MSCOs—often misinterpreted as a phenomenon unique to general relativity—emerge under Newtonian conditions due to extended mass effects and density inhomogeneity. Meanwhile, under prolate deformation, MSCOs are not formed.
[…]
As another astrophysical application, we consider the effect of dark matter on the orbit of the S2 star around the Galactic Center. The S2 star has an orbital semi-major axis of $a_{\mathrm{S2}}=970\mathrm{\,au}\approx 1.45\times 10^{16}\mathrm{\,cm}$ and eccentricity $e_{\mathrm{S2}}=0.88$, and orbits the supermassive object at the center of the Milky Way, with mass $M_{\bullet}=4.3\times 10^{6}M_{\odot}$ (GRAVITY Collaboration et al., 2020). Although our quasi-circular approximation does not strictly apply to such a highly eccentric orbit, we can still estimate the order of magnitude of the effect by considering the mean enclosed density. The mean mass density enclosed within $a_{\mathrm{S2}}$, set by the central mass, is given by $\bar{\rho}_{\mathrm{S2}}=\frac{3M_{\bullet}}{4\pi a_{\mathrm{S2}}^{3}}\approx 6.7\times 10^{-10}\mathrm{\,g/cm^{3}}$. For illustration, if we assume that the maximum allowed amount of dark matter (corresponding to a mass fraction of 0.03% or $1200M_{\odot}$ (GRAVITY Collaboration et al., 2020)) is uniformly distributed within the S2 orbital radius, the resulting density is $\rho_{\mathrm{DM}}=1.9\times 10^{-13}\mathrm{\,g/cm^{3}}$. This yields a periapsis shift correction of $\Delta\phi_{\mathrm{p}}\approx-2.6\times 10^{-3}\mathrm{\,rad/orbit}$, which is comparable to the general relativistic contribution ($3.5\times 10^{-3}\mathrm{\,rad/orbit}$). The correction due to the deformation effect is expected to be of a similar order, or possibly smaller, depending on the details of the deformation parameter. This implies that, if a significant amount of deformation is present in the mass distribution near the Galactic Center, its impact on the periapsis shift could in principle compete with the relativistic effect. More realistically, in standard scenarios, such as those based on the NFW halo profile (see e.g., (Gnedin & Primack, 2004)), the dark matter density at the Galactic center is expected to be $O(10^{-19})\mathrm{\,g/cm^{3}}$. However, if the formation of a dark matter spike due to adiabatic growth of the central black hole is considered, the density can reach much higher values, potentially up to $O(10^{-16})\mathrm{\,g/cm^{3}}$ or even greater (Gondolo & Silk, 1999; Lacroix, 2018; Sadeghian, Ferrer & Will, 2013), in which case the extended mass effect (i.e., the spherically symmetric component) and the associated deformation effect (the non-spherical correction), would become more pronounced due to the localization of the dark matter density.
@@
We will report the results of our post-Newtonian analysis in a separate paper, which examines the competition between the general-relativistic effects and the combined extended mass and deformation effects on the periapsis shift of quasi-circular orbits in a system featuring a black hole at the center of a spheroidal mass distribution. Furthermore, the extended mass and deformation effects on the periapsis shift of quasi-circular orbits in static, axisymmetric spacetimes warrant future work (Iorio, 2024). In addition, generalizing the general formula for the periapsis shift of quasi-circular orbits in static, spherically symmetric systems (2) to static, axisymmetric spacetimes is a promising topic for future research.
Acknowledgements
@@
The author gratefully acknowledges the useful comments and discussions provided by Kota Ogasawara, Tomohiro Harada, Masashi Kimura, Keiju Murata, Takeshi Chiba, Masaya Amo, Akihito Katsumata, Takashi Mishima, Yohsuke Takamori, Takaaki Ishii, Hideki Ishihara, Ken-ichi Nakao, Yasufumit Kojima, Masaaki Takahashi, Rohta Takahashi, Shinji Koide, Hirotaka Yoshino, and So Aina. The author would like to thank the anonymous referee for their constructive comments, which improved the quality of this manuscript. This work was supported by JSPS KAKENHI Grant Nos. JP22K03611, JP23KK0048, and JP24H00183.
DATA AVAILABILITY
No data were created or analyzed in this article.
@@ A.1 Oblate case ($a>b$)
$C_{5,0}={}$ $\frac{63(1-e^{2})}{128e^{10}}\bigg{(}\frac{\arcsin e}{e\sqrt{1-e^{2}}}-1-\frac{2e^{2}}{3}-\frac{8e^{4}}{15}-\frac{16e^{6}}{35}$
$\quad-\frac{128e^{8}}{315}\bigg{)}.$ (74)
@@
$C_{3,1}={}$ $\frac{35(1-e^{2})}{8e^{8}}\bigg{(}1-\frac{e^{2}}{3}-\frac{2e^{4}}{15}-\frac{8e^{6}}{105}$
$\quad-\frac{\sqrt{1-e^{2}}}{e}\arcsin e\bigg{)},$ (78)
$C_{4,1}={}$ $\frac{315(1-e^{2})}{64e^{10}}\bigg{(}1-\frac{e^{2}}{3}-\frac{2e^{4}}{15}-\frac{8e^{6}}{105}-\frac{16e^{8}}{315}$
$-\frac{\sqrt{1-e^{2}}}{e}\arcsin e\bigg{)}.$ (79)
@@ A.2 Prolate case ($a<b$)
$C_{4,0}={}$ $\frac{35}{64e^{8}}\bigg{(}-1+\frac{11e^{2}}{3}-\frac{73e^{4}}{15}+\frac{93e^{6}}{35}$
$\quad+\frac{(1-e^{2})^{4}}{e}\operatorname{arctanh}e\bigg{)},$ (84)
$C_{5,0}={}$ $\frac{63}{128e^{10}}\bigg{(}1-\frac{14e^{2}}{3}+\frac{128e^{4}}{15}-\frac{158e^{6}}{21}+\frac{193e^{8}}{63}$
$\quad-\frac{(1-e^{2})^{5}}{e}\operatorname{arctanh}e\bigg{)},$ (85)
@@
$C_{4,1}={}$ $\frac{315}{64e^{10}}\bigg{(}-1+\frac{11e^{2}}{3}-\frac{73e^{4}}{15}+\frac{93e^{6}}{35}-\frac{138e^{8}}{315}$
$\quad+\frac{(1-e^{2})^{4}}{e}\operatorname{arctanh}e\bigg{)}.$ (90)